40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 74 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  13810   Thu May 3 10:40:43 2018 johannesConfigurationGeneralAS port laser injection

Instead of trying to couple the fiber output into the interferometer, I'm doing the reverse and maximize the amount of interferometer light going into the fiber. I set up the mode-matching solution shown in attachment #1 and started tweaking the lens positions. Attachment #2 shows the setup on the AS table. After the initial placement I kept moving the lenses in the green arrow directions and got more and more light into the fiber.

When I stopped this work yesterday I measured 86% of the AS port light coming out the other fiber end, and I have not yet reached a turning point with moving the lenses, so it's possible I can tickle out a little more than that.

It occured to me though that I may have been a little hasty with the placement of the mirror that in attachment #2 redirects the beam which would ordinarily go to AS55. For my arm ringdown measurements this doesn't matter, I could actually place it even before the 50/50 beamsplitter that sends light onto AS110 and double the amount of light going into the IFO. What signals are needed for the Guoy phase measurement? Is AS 110 sufficient, or do we need AS55?

Attachment 1: mm_solution_AStable.png
mm_solution_AStable.png
Attachment 2: AStable_beampath.pdf
AStable_beampath.pdf
  13809   Thu May 3 09:56:42 2018 SteveHowToSEIpreparation of load cell measurement at ETMX

[ Dennis Coyne'  precision answer ]

Differential Height between Isolators

According to a note on the bellows drawing (D990577-x0/A), the design life of the bellows at ± 20 minutes rotational stroke is 10,000 cycles. A 20 minute angular (torsional) rotation of the bellows corresponds to 0.186" differential height change across the 32" span between the chamber support beams (see isolator bracket, D000187-x0/B).

Another consideration regarding the bellows is the lateral shear stress introduced by the vertical translation. The notes on the bellows drawing do not give lateral shear limits. According to MDC's web page for formed bellows in this size range the lateral deflection limit is approximately 10% of the "live length" (aka "active length", or length of the convoluted section). According to the bellows drawing the active length is 3.5", so the maximum allowable lateral deflection should be ~0.35".

Of course when imposing a differential height change both torsional and lateral shear is introduced at the same time. Considering both limits together, the maximum differential height change should be < 0.12".

One final consideration is the initial stress to which the bellows are currently subjected due to a non-centered support beam from tolerances in the assembly and initial installation. Although we do not know this de-centering, we can guess that it may be of the order of ~ 0.04". So the final allowable differential height adjustment from the perspective of bellows stress is < 0.08".   Steve:  accumulated initial stress is unknown.  We used to adjust the original jack screws for IFO aligment in the early days of ~1999. This kind of adjustment was stopped when we realized how dangereous it can be. The fact is that there must be unknown amount of accumulated initial stress. This is my main worry but I'm confident that 0.020" change is safe.

So, with regard to bellows stress alone, your procedure to limit the differential height change to <0.020" is safe and prudent.

However, a more stringent consideration is the coplanarity requirement (TMC Stacis 2000 User's Manual, Doc. No. SERV 04-98-1, May 6, 1991, Rev. 1), section 2, "Installation",which stipulates < 0.010"/ft, or < 0.027" differential height across the 32" span between the chamber support beams. Again, your procedure to limit the differential height change to < 0.02" is safe.

Centered Load on the STACIS Isolators

According to the TMC Stacis 2000 User's Manual (Document No. SERV 04-98-1, May 6, 1991, Rev. 1), section 2, "Installation", typical installations (Figure 2-3) are with one payload interface plate which spans the entire set of 3 or 4 STACIS actuators. Our payload interface is unique.

Section 2.3.1, "Installation Steps": "5. Verify that the top of each isolator is fully under the payload/interface plate; this is essential to ensure proper support and leveling. The payload or interface plate should cover the entire top surface of the Isolator or the entire contact area of the optional jack."

section 2.3.2, "Payload/STACIS Interface": "... or if the supporting points do not completely cover the top surface of each Isolator, an interface plate will be needed."

The sketch in Figure 2-2 indicates an optional leveling jack which appears to have a larger contact surface area than the jacks currently installed in the 40m Lab. Of course this is just a non-dimensioned sketch. Are the jacks used by the 40m Lab provided by TMC, or did we (LIGO) choose them? I beleive Larry Jones purchased them.

A load centering requirement is not explicitly stated, but I think the stipulation to cover the entire top surface of each actuator is not so much to reduce the contact stress but to entire a centered load so that the PZT stack does not have a reaction moment.

From one of the photos in the 40m elog entry (specifically jack_screw.jpg), it appears that at least some isolators have the load off center. You should use this measurement of the load as an opportunity to re-center the loads on the Isolators.

In section 2.3.3, "Earthquake Restraints" restraints are suggested to prevent damage from earth tremors. Does the 40m Lab have EQ restraints? Yes, it has

Screw Jack Location

I could not tell where all of the screw jacks will be placed from the sketch included in the 40m elog entry which outlines the proposed procedure.

Load Cell Locations

The sketch indicates that the load cells will be placed on the center of the tops of the Isolators. This is good. However while discussing the procedure with Gautam he said that he was under the impression that the load cell woudl be placed next to the leveling jack, off-center. This condition may damage the PZT stack. I suggest that the leveling jack be removed and replaced (temporarily) with the load cell, plus any spacer required to make up the height difference. Yes

If you have any further question, just let me know.

    Dennis

 

 

Dennis Coyne
Chief Engineer, LIGO Laboratory
California Institute of Technology
MC 100-36, 1200 E. California Blvd.

 

 

 

  13808   Thu May 3 00:42:38 2018 KevinUpdatePonderSqueezeCoil driver contribution to squeezing noise budget

In light of the discussion at today's meeting, Guantanamo and I looked at how the series resistance for the test mass coil drivers limits the amount of squeezing we could detect.

The parameters used for the following calculations are:

  • 4.5 kΩ series resistance for the ETM's (this was 10 kΩ in the previous calculations, so these numbers are a bit worse); 15 kΩ for the ITM's
  • 100 ppm transmissivity on the folding mirrors giving a PRC gain of 40
  • PD quantum efficiency of 0.88

Since we need to operate very close to signal recycling, instead of the current signal extraction setup, we will need to change the macroscopic length of the SRC. This will change the mode matching requirements such that the current SRM does not have the correct radius of curvature. One solution is to use the spare PRM which has the correct radius of curvature but a transmissivity of 0.05 instead of 0.1. So using this spare PRM for the SRM and changing the length of the SRC to be the same as the PRC we can get

  • 0.63 dBvac of squeezing at 205 Hz for 1 W incident on the back of PRM
  • 1.12 dBvac of squeezing at 255 Hz for 5 W incident on the back of PRM

This lower transmissivity for the SRM also reduces the achievable squeezing from the current transmissivity of 0.1. For an SRM with a transmissivity of 0.15 (which is roughly the optimal) we can get

  • 1 dBvac of squeezing at 205 Hz for 1 W incident on the back of PRM
  • 1.7 dBvac of squeezing at 255 Hz for 5 W incident on the back of PRM

The minimum achievable squeezing moves up from around 205 Hz at 1 W to 255 Hz at 5 W because the extra power increases the radiation pressure at lower frequencies.

  13807   Wed May 2 21:39:33 2018 gautamConfigurationALSIR ALS for EY

The new K6XS mounts I ordered have arrived. I want to install one of them at the Y-end. I can't find a picture of the current layout but it exists as there is a hardcopy affixed to the ETMY chamber door, Steve, can we dig this up and put it in the wiki? In any case, the current beam going into the fiber is the pickoff from the post-SHG harmonic separator. I'd like to change the layout a bit, and use a pickoff before the doubling oven, but looking at the optical table, this seems like a pretty involved task and would probably require large scale optical hardware rearrangement. In any case, the MM of the green beam into the Y-arm is <50%, so I would like to redo that as well. Does anyone know of a measurement of the mode from the Lightwave NPRO that is installed at EY? I think Annalisa is the one who installed this stuff, but I can't find an actual NPRO mode measurement in her elog thread.


Found it: elog4874, elog8436. I updated the laser inventory page to link the lasers in use to the most recent mode measurements I could find on the elog. I guess ideally we should also link the AM/PM response measurements.

------------------------------------------------------------------------------------------------------

SV  ETMY optical table layout  

     as of 3-31-2016

The oplev path was optimized with AR coated lenses and new He/Ne laser Jan 24, 2017

  13806   Wed May 2 10:03:58 2018 SteveHowToSEIpreparation of load cell measurement at ETMX

Gautam and Steve,

We have calibrated the load  cells. The support beams height monitoring is almost ready.

The danger of this measurment that  the beams height changes can put shear and torsional forces on this formed (thin walled) bellow

They are designed for mainly axial motion.

The plan is to limit height change to 0.020" max

0, center oplev at X arm locked

1, check that  jack screws are carrying full loads and set height indicator dials to zero ( meaning: Stacis is bypassed )

2, raise beam height with aux leveling wedge  by 0.010"  on all 3 support point and than raise it an other 0.005"

3, replace levelling wedge with load cell that is centered and shimmed.     Dennis   Coyne pointed out that the Stacis foot has to be loaded at the center of the foot and formed bellow can shear at their limits.

4, lower the support beam by 0.005" ......now full load on the cells

Note: jack screw heights will not be adjusted or  touched.......so the present condition will be recovered

Quote:

We could use similar load cells   to make the actual weight measurement on the Stacis legs. This seems practical in our case.

I have had bad experience with pneumatic Barry isolators.

Our approximate max compression loads are 1500 lbs on 2 feet and 2500 lbs on the 3rd one.

 

 

Attachment 1: loadcellCAL500.pdf
loadcellCAL500.pdf
Attachment 2: 3loadcellwcontr.jpg
3loadcellwcontr.jpg
Attachment 3: loadcellLocation.pdf
loadcellLocation.pdf
Attachment 4: DSC01009.JPG
DSC01009.JPG
Attachment 5: jack_screw.jpg
jack_screw.jpg
Attachment 6: ETMX_NW_foot_STACIS.pdf
ETMX_NW_foot_STACIS.pdf
  13805   Tue May 1 19:37:50 2018 gautamUpdateGeneralDARM actuation estimate

Here is an updated plot - the main difference is that I have added a trace that is the frequency domain wiener filter subtraction of the coherent power between the L_X and L_Y time series. I tried reproducing the calculation with the time domain wiener filter subtraction as well, using half of the time series (i.e. 5 mins) to train the wiener filter (with L_X as target and L_Y as witness), but I don't get any subtraction above 5 Hz on the half of the data that is a test data set. Probably I am not doing the pre-filtering correctly - I downsampled the signal to 1 kHz, weighted it by low passing the signal above 40 Hz and trained the Wiener filter on the resulting time series. But this frequency domain Wiener filter subtraction should be at least a lower bound on DARM - indeed, it is slightly lower everywhere than simply taking the time domain subtraction of the two data streams.

To do:

  • Re-measure calibration numbers used.
  • Redo calculation once the numbers have been verified.

Putting a slightly cleaned up version of this plot in now - I'm only including the coherence-inferred DARM estimate now instead of the straight up time domain subtraction. So this is likely to be an underestimate. At low (<10 Hz) frequencies, the time domain computation lines up fairly well, but I suspect that I am getting huge amounts of spectral leakage (see Attachment #2) in the way I compute the spectrum using scipy's filtering routine (once the Wiener filter has been computed). Note that Attachment #2, I didn't break up the data into a training/testing set as in this case, we just care about the one-off offline performance in order to get an estimate of DARM.

The python version of the wiener filter generating code only supports [b,a] output of the digital filter, an sos filter might give better results. Need to figure out the least painful way of implementing the low-noise digital filtering in python...

Attachment 1: darmEst.pdf
darmEst.pdf
Attachment 2: darmEst_time.pdf
darmEst_time.pdf
  13804   Tue May 1 15:23:18 2018 KiraUpdatePEMnew ADC channel setup issue

[Kira, Johannes]

I connected up the channels for the ADC Acromag a while back and we were planning to install it today so that we could set up a new channel for the out of loop sensor. Unfortunately, the Acromag seems to be broken. We connected up a precision 10V voltage to one of the channels, but the Acromag read out ~7V and it kept fluctuating. Even after calibration, we still got the same result. When enabling the legacy support, we got ~11V. But when we measured the voltage at the screw terminals with a multimeter and it showed 10V, so the issue is not with the wiring. All of the channels have this same issue. We will be ordering more Acromags soon, so hopefully we'll be able to set up the channel soon. I've attached a picture of the Acromag along with the front panel with the channels labeled

Attachment 1: IMG_20180501_152014.jpg
IMG_20180501_152014.jpg
  13803   Tue May 1 11:15:19 2018 KiraUpdatePEMPID Quixote

I added an out of loop sensor to the can by placing the lab temperature sensor inside the can. I'm not sure which channel is logging this temperature though. I also noticed that the StripTool still had the old misspelled name for the temperature readout so I fixed that as well.

I've attached a picture of the setup.

Quote:

Increased the Integral gain (from -1 to -4) on the EX temperature controller. This didn't work a few weeks ago, but now with the added P gain, it seems stable. Daily temperature swings are now ~3x smaller.

Notes for Kira on what we need to do tomorrow (Friday):

  1. add the MEDM screen EPICS values to the DAQ so that we can plot those trends DONE
  2. add the out-of-loop sensor to the EX can
  3. reboot the AUX-EX so we can pick up the new channels and the fixed spelling of the old channels DONE
  4. Re-install EX seismometer and hook up seismometer channels to PEM DAQ so we can start testing its performance.

For those who are flabbergasted by the way I calibrated the TEMP_MON channel from volts to deg C, here's how:

XMgrace->Data->Transformations->Geometric Transforms...

use the 'scale' and 'translate' fields to change the slope and offset for calibration in the obvious ways

 

Attachment 1: IMG_20180501_154826.jpg
IMG_20180501_154826.jpg
  13802   Tue May 1 08:04:13 2018 Jon RichardsonConfigurationElectronicsPSL-Aux. Laser Phase-Locked Loop

[Jon, Gautam, Johannes]

Summary: In support of making a proof-of-concept RF measurement of the SRC Gouy phase, we've implemented a PLL of the aux. 700mW NPRO laser frequency to the PSL. The lock was demonstrated to hold for minutes time scales, at which point the slow (currently uncontrolled) thermal drift of the aux. laser appears to exceed the PZT dynamic range. New (temporary) hardware is set up on an analyzer cart beside the PSL launch table.

Next steps:

- Characterize PLL stability and noise performance (transfer functions).

- Align and mode-match aux. beam from the AS table into the interferometer.

- With the IFO locked in a signal-recycled Michelson configuration, inject broadband (swept) AM sidebands via the aux. laser AOM. Coherently measure the reflection of the driven AM from the SRC.

- Experiment with methods of creating higher-order modes (partially occluding the beam vs. misaligning into, e.g., the output Faraday isolator). The goal is identify a viable techinque that is also possible at the sites, where the squeezer laser serves as the aux. laser.

The full measurement idea is sketched in the attached PDF.

IMG_2551.jpg
PSL-Aux. beat note sensor on the PSL launch table.
IMG_2552.jpg
Feedback signal to aux. laser PZT.
IMG_2553.jpg
PLL electronics cart.

 

Attachment 1: IMG_2553.jpg
IMG_2553.jpg
Attachment 4: src_gouy_phase_v3.pdf
src_gouy_phase_v3.pdf src_gouy_phase_v3.pdf src_gouy_phase_v3.pdf src_gouy_phase_v3.pdf
  13801   Mon Apr 30 23:13:12 2018 KevinUpdateComputer Scripts / ProgramsDataViewer leapseconds

I was trying to plot trends (min, 10 min, and hour) in DataViewer and got the following error message

Connecting.... done
 mjd = 58235
leapsecs_read()
  Opening leapsecs.dat
  Open of leapsecs.dat failed
leapsecs_read() returning 0
frameMemRead - gpstimest = 1208844718

 

thoough the plots showed up fine after. Do we need to fix something with the leapsecs.dat file?

  13800   Mon Apr 30 15:36:18 2018 KiraUpdatePEMfinal setup sketch

I've attached a sketch of how the panel will be mounted. We should make a small rectangular box that would raise the panel from the block by 1 cm or so to allow the cables to fit into the hole in the block without getting bent. It also has to be airtight so maybe having a thin layer of rubber between the mount and block would be good.

Attachment 1: mount.png
mount.png
  13799   Sun Apr 29 22:53:06 2018 gautamUpdateGeneralDARM actuation estimate

Motivation:

We'd like to know how much actuation is required on the ETMs to lock the DARM degree of freedom. The "disturbance" we are trying to cancel is the seismic driven length fluctuation of the arm cavity. In order to try and estimate what the actuation required will be, we can use data from POX/POY locks. I'd collected some data on Friday which I looked at today. Here are the results.

Method:

  • I collected the error and control signals for both arm cavities while they were locked to the PSL.
  • Knowing the POX/POY sensing response and the actuator transfer functions, we can back out the free running displacements of the two arm cavities.
    • I used numbers from the cal filters which may not be accurate (although POX sensing response which was recently measured).
    • But the spectra computed using this method seem reasonable, and the X and Y arm asds line up around 1 Hz (albeit on a log scale).
    • In this context, L_X is really a proxy for |f_X - f_{MC}| and similarly for L_Y so I think the algebra works out correctly.
    • I didn't include any of the violin mode/AA/AI filters in this calculation.
  • Having calculated the arm cavity displacements, I computed "DARM" as L_y- L_x and then plotted its asd.
  • For good measure, I also added the quadrature sum of 4 optics' displacement noise as per the 40m GWINC model - there seems to be a pretty large discrepancy, not sure why.

If this approach looks legit, I will compute the control signal that is required to stabilize this level of disturbance using the DARM control loop, and see what is the maximum permissible series resistance we can use in order to realize this stabilization. We can then compare various scenarios like different whitening schemes, with/without Barry puck etc, and look at coil driver noise levels for each of them. 

Attachment 1: darmEst.pdf
darmEst.pdf
  13798   Fri Apr 27 18:42:02 2018 ranaUpdatePEMnew Seis temp chans

for whatever reason, I am unable to get minute or second trends from nodus for any channels (IMC, PEM, etc) since the reboot. has there been some more recent FB failure or is this still a bug since last years FB catastrophe?

  13797   Fri Apr 27 16:55:31 2018 gautamUpdateGeneralEY area access blocked

Steve was calibrating the load cells at the EY table with the crane - we didn't get through the full procedure today, so the area near the EY table is kind of obstructed. The 100kg donut is resting on the floor on the North side of the EY table and is still connected to the crane. There are stopper plates underneath the donut, and it is still connected to the crane. Steve has placed cones around the area too. The crane has been turned off.

  13796   Fri Apr 27 01:36:02 2018 gautamConfigurationALSIR ALS noise performance

Summary:

My goal was to do some further characterization of the IR ALS system tonight. With POX as an OOL sensor, I measured an RMS displacement noise of  8 pm with the arm under ALS control. I calculated the CARM linewidth to be 77 Hz (=10.3 pm) for the 40m parameters, assuming 30ppm arm loss. Fuurthermore, this number is 3x better than the 24 pm RMS quoted in the Izumi et. al. paper. Of course I am quoting the best results from my efforts tonight. Conclusions:

  1. [Attachment #1] --- With XARM locked using POX, the ALS beat noise (i.e. Phase Tracker output noise) lines up well with the reference we have been using for some time now (and indeed, is better in some places).
  2. [Attachment #2] --- With the arm locked on ALS and POX as an OOL sensor, I measured performance comparable to this measurement we did sometime last year. Anomalies in this measurement and the one above were what precipitated the IMC noise investigation.
  3. [Attachment #3] --- The above two attachments are not the whole story. During the day, I get significantly worse performance (so much so that I couldn't even do the handoff to ALS control). But in 5 minutes of measurement, the ALS noise seems quite stationary.
  4. [Attachment #4] --- This is really the same as Attachment #2, but I wanted to overlay some vlines. Maybe this is a clue to some 60 Hz / ground loop issues, but the RMS has significant contribution from these harmonics. Tmrw, I will add the old measurement overlaid to this plot (and for what its worth, the Izumi et. al. spectrum as well).
  5. [Attachment #5] --- With the arm under ALS control, I was able to maintain the lock for a solid hour (and more as I write up this elog). Somehow inkscape screwed up the fonts, but main point here is that TRX is stable to within 10% throughout the observation time.

Since the stability and noise seemed quite good, I decided to collect some arm scan data to give to our modeSpec SURFs to practice fitting (which is the short dip in TRX in Attachment #4). Although after the discussion with Rana today, I think it may be that we want to do this measurement in reflection and not transmission, and look for a zero crossing in the PDH signal. In any case, I was able to scan 7 FSRs without any issues. I will upload the data to some git repo. GPS start time is 1208850775, sweep was 3mins long.

I think the next step here is to noise-budget this curve. At least the DFD noises

Attachment 1: 2018_04_BeatMouth_POX.pdf
2018_04_BeatMouth_POX.pdf
Attachment 2: 2018_04_BeatMouth.pdf
2018_04_BeatMouth.pdf
Attachment 3: ALSSpecgram.pdf
ALSSpecgram.pdf
Attachment 4: ALS_ASD.pdf
ALS_ASD.pdf
Attachment 5: ALSstab.pdf
ALSstab.pdf
  13795   Thu Apr 26 23:00:42 2018 ranaUpdatePEMnew Seis temp chans

After fixing the spelling of the EX temperature readback, I also added all of the MEDM sliders to the C0EDCU.ini file (making sure to add an even number of channels). Restarted FB (after installing telnet on rossa):

telnet fb 8083

> shutdown


preferred method of posting DataViewer images: print as a SVG image (since its vectorized). Then from the command line do:

inkscape steven.svg --export-pdf=vass.pdf

Attachment 1: chans.pdf
chans.pdf
  13794   Thu Apr 26 20:22:21 2018 KevinUpdatePEMADC common mode rejection with new seismometer connections

Yesterday I wired the outputs from the seismometers directly to the ADC input bypassing the old AA board circuit as is described in this elog. The old circuit converted the single-ended output from the seismometers to a differential signal. Today I looked at whether 60 Hz noise is worse going directly into the ADC due to the loss of the common mode rejection previously provided by the conversion to differential signals.

I split the output from the BS Z seismometer to the new board and to an SR785. On the SR785 I measured the difference between the inner and outer conductors of the seismometer output, i.e. A-B with A the center conductor and B the outer conductor, with grounded input. At the same time I took a DTT spectrum of C1:PEM-SEIS_BS_Z_IN1. Both spectra were taken with 1 Hz bandwidth and 25 averages. The setup is shown in attachment 1.

The spectra are shown in attachment 2. The DTT spectrum was converted from counts to volts by multiplying by 2 * 10 V/32768 cts where the extra factor of 2 is from converting from single-ended to differential input. If there was common 60 Hz noise that the ADC was picking up we would expect to see less noise at 60 Hz in the SR785 spectrum measured directly at the output from the seismometer since that was a differential measurement. Since both spectra have the same 60 Hz noise, this noise is differential.

Attachment 1: setup.pdf
setup.pdf
Attachment 2: seismometerASD.pdf
seismometerASD.pdf
  13793   Thu Apr 26 19:46:26 2018 ranaUpdatePEMPID Quixote

Increased the Integral gain (from -1 to -4) on the EX temperature controller. This didn't work a few weeks ago, but now with the added P gain, it seems stable. Daily temperature swings are now ~3x smaller.

Notes for Kira on what we need to do tomorrow (Friday):

  1. add the MEDM screen EPICS values to the DAQ so that we can plot those trends DONE
  2. add the out-of-loop sensor to the EX can
  3. reboot the AUX-EX so we can pick up the new channels and the fixed spelling of the old channels DONE
  4. Re-install EX seismometer and hook up seismometer channels to PEM DAQ so we can start testing its performance.

For those who are flabbergasted by the way I calibrated the TEMP_MON channel from volts to deg C, here's how:

XMgrace->Data->Transformations->Geometric Transforms...

use the 'scale' and 'translate' fields to change the slope and offset for calibration in the obvious ways

Attachment 1: dv.pdf
dv.pdf
  13792   Thu Apr 26 18:58:21 2018 BruceConfigurationALSNew look EX Fiber coupling - pol stability

  13791   Thu Apr 26 11:24:50 2018 gautamConfigurationALSNew look EX Fiber coupling - pol stability

Here is a first look at the overnight stability. For the temperature, I used the calibration I found in the old psl database file, seems to give sensible results. It's only 15 hours of data plotted, so we don't see the full 24 hour temperature swing, but I think it is safe to say that for the EX fiber, the dominant cause of the "waveplate effect" is not in fact temperature drift. The polarization extinction is still better than 10dB in the entire period of observation though... I'm going to push ahead with a beat spectrum measurement, though there is room for improvement in the input coupling alignment to fiber special axes.

The apparent increase in these plots towards the end of the 15 hour period is because the lights on the PSL table were switched on.


Annoyingly, it seems like the PSL NPRO channels (which I have hijacked to do this test) do not have minute trend data directly accessible from NDS2. Not sure whether this is an NDS2 problem, or something missing in the way the channels are setup with Acromag. Probably the former, as I am able to generate minute trend plots with dataviewer. I forget whether this is the same as the infamous minute trend problem. Second trend data is available though, and is what I used to make these plots...

Attachment 1: polStab.pdf
polStab.pdf
  13790   Thu Apr 26 09:35:49 2018 KevinUpdatePEMPEM Anti-Alias wiring

I wired all 32 channels going to the AA board directly to the ADC as described in the previous log. However, instead of using the old AA board and bypassing the whole circuit, I just used a breakout board as is shown in the first attachment. I put the board back in the rack and reconnected all of the cables.

The seismic BLRMs appear to be working again. A PSD of the BS seismometers is shown in attachment 2. Tomorrow I'll look at how much the ADC alone is suppressing the common mode 60 Hz noise on each of the channels.

Steve: 5 of ADC DAC In Line Test Boards [ D060124 ] ordered. They should be here within 10 days.

Attachment 1: board.jpg
board.jpg
Attachment 2: SeismometerPSD.pdf
SeismometerPSD.pdf
  13789   Wed Apr 25 19:09:37 2018 gautamConfigurationALSNew look EX Fiber coupling

Summary:

I implemented most of the things outlined in my previous elog. Implementing the a la mode solution after including all lenses, I managed to achieve >90% mode-matching into the fiber. Power monitor PD has not been re-installed yet, neither has the bracket I removed. The polarization monitoring setup on the PSL table has now been hooked up to the EX fiber, let's see how it does overnight. All quoted power measurements in this elog were made with the Ophir power meter (filter off).

Details:  

Attachment #1 shows the implemented MM solution. I did not include the PBS substrate in the calculation, maybe that will help a little.

Attachment #2 shows the new layout. The beam is a little low on the PBS and HWP - I will swap these out to mounts with slightly lower height, that should improve the situation a little. There is no evidence of clipping, and the beam clears all edges by at least 3 beam diameters.

Attachments #3 and #4 show the EX fiber before and after cleaning respectively. Seems like the cleaning was successful.

Attachment #5 shows the beam incident on the coupler with on an IR card. This beam only goes through a QWP, lens, BS and 45 degree steering mirror, so I'm not sure what's responsible for the large halo around the main beam. There is significant power in the halo too - I measured 25mW right before the coupler, but if I use an iris to try and cut off the halo, the power is measured to be ~19mW.

Alignment Procedure:

  • Connect spare fiber such that I can monitor coupled power (minus fiber losses and joint loss) at EX table.
  • Use Fluke fault analyzer to align input and collimator modes coarsely.
  • Monitored coupled power continuously using Fiber Power Meter (although MM calculations were made with Ophir, this was more convenient for "Live" viewing).
  • Tweaked one available steering mirror and K6XS axes to maximize coupled power. 
  • Tweaked lens positions slightly to see if significant improvement could be made.
  • After optimizing, I measured 17.1mW coming out of the EX fiber at the PSL table. As mentioned earlier, the input power is tricky to measure given the large amount of junk light around the main mode. But I measured 18.6 mW after the iris. So this is ~95%. In any case, safe to say that we are waaaay better than the previous situation of 380uW out of 1.9mW. 
  • Added PBS and HWP to cut the incident power to 1.6mW. I measured 1.2mW on the PSL table. Probably adding the PBS screwed up the MM a bit, to be tweaked tomorrow. 
  • I had moved the Green shutter a bit during this work - as a result, the Green REFL was not making it back to the REFL PD. I remedied this, and EX Green TEM00 mode was locked to the arm. GTRX of ~0.4 was recovered, which is around the number I'm used to seeing.
Attachment 1: EX_fiber_MM.pdf
EX_fiber_MM.pdf
Attachment 2: IMG_6977.JPG
IMG_6977.JPG
Attachment 3: IMG_6972.JPG
IMG_6972.JPG
Attachment 4: IMG_6974.JPG
IMG_6974.JPG
Attachment 5: IMG_6976.JPG
IMG_6976.JPG
  13788   Wed Apr 25 17:44:39 2018 ArnoldUpdatePEMPEM Anti-Alias wiring

Related image

 

  13787   Tue Apr 24 21:19:08 2018 KevinUpdatePEMSeismometer BLRMs

In the ongoing attempt to recover the seismometer BLRMS, I removed the AA board from the rack and modified the BS seismometer Z channel. The BS_Z BLRMs seem to be recovered after this modification.

I removed the three resistors from the output of the circuit and wired the input and from the seismometer directly to the input to the ADC. The modified schematic is shown in attachment 1. Attachments 2 and 3 show the top and bottom of the modified board. The board is doing nothing now other than serving as a connector for this channel.

I put the board back in the rack and injected a 2 Vpp signal into the BS_Z channel and saw +/- 1600 cts in C1PEM-SEIS_BS_Z. I then plugged the seismometer back into the board and took the spectrum shown in attachment 4. This shows the working Z channel giving a reasonable seismic spectrum. Note that X and Y are not modified yet.

If there are no objections, I will modify all the other channels on the board in the same way tomorrow.

Attachment 1: modified_schematic.pdf
modified_schematic.pdf
Attachment 2: top.jpg
top.jpg
Attachment 3: bottom.jpg
bottom.jpg
Attachment 4: BS_Seis_PSD.pdf
BS_Seis_PSD.pdf
  13786   Tue Apr 24 18:54:15 2018 gautamConfigurationALSProposed changes to EX fiber coupling

I started working on the EX table. Work is ongoing so I will finish this up later in the evening, but in case anyone is wondering why there is no green light...

  1. EX laser shutter was closed.
  2. Disconnected EX input to the beat mouth at the PSL table in order to avoid accidentally frying the PDs.
  3. Prepared new optomechanics hardware
    • To my surprise, I found a bubble-wrapped K6XS mount (the one with locking screws for all DoFs) on the SP table. No idea where this came from or who brought it here, or how long it has been here, but I decided to use it nevertheless.
    • Prepared f = 200mm and f = -200mm lenses on traveling mounts (Thorlabs DT12, lenses are also Thorlabs, AR1064).
    • Made a slight translation of the beam path towards the north to facilitate going through the center of the mounted lenses.
    • Temporarily removed a beam dump from next to the final steering mirror before the Green REFL PD, and also removed one of the brackets between the enclosure and the table for ease of laying out components. These will be replaced later.
  4. Installed this hardware on the PSL table, roughly aligned beam path.
    • Beam now goes through the center of all lenses and is hitting the collimator roughly in the center.

To do in the eve:

  1. Clean fiber and connect it to the collimator.
  2. Optimize mode-matching as best as possible.
  3. Attenuate power using PBS and HWP so as to not damage the BeatMouth PD (Pthresh = 2mW). These are also required to make the polarizations of the EX coupled light (S-pol) and PSL (P-pol) go along the same axis of the PM fiber.
  4. Re-install temporarily removed beam dump and bracket on EX table.
  5. Re-install EX power monitoring PD.
  6. Measure beat frequency spectrum.
Quote:

Motivation: I want to make another measurement of the out-of-loop ALS beat noise, with improved MM into both the PSL and EX fibers and also better polarization control. For this, I want to make a few changes at the EX table. 

Barring objections, I will start working on these changes later today.


gautam 1245am: Fiber cleaning was done - I'll upload pics tomorrow, but it seems like the fiber was in need of a good cleaning. I did some initial mode-matching attempts, but peaked at 10% MM. Koji suggested not going for the final precisely tunable lens mounting solution while trying to perfect the MM. So I'll use easier to move mounts for the initial tuning and then swap out the DT12s once I have achieved good MM. Note that without any attenuation optics in place, 24.81mW of power is incident on the collimator. In order to facilitate easy debugging, I have connected the spare fiber from PSL to EX at the PSL table to the main EX fiber - this allows me to continuously monitor the power coupled into the fiber at the EX table while I tweak lens positions and alignment. After a bit of struggle, I noticed I had neglected a f=150mm lens in my earlier calculation - I've now included it again, and happily, there seems to be a solution which yields the theoretical 100% MM efficiency. I'll work on implementing this tomorrow..

  13785   Tue Apr 24 15:59:23 2018 SteveUpdateWiki AP table layout 20180328
Quote:

ETMX table layout uploaded with beam paths to the wiki.   laugh

The pdf file is uploaded into the wiki.

Attachment 1: DSC00668AP20180328.png
DSC00668AP20180328.png
  13784   Tue Apr 24 11:31:59 2018 gautamConfigurationALSProposed changes to EX fiber coupling

Motivation: I want to make another measurement of the out-of-loop ALS beat noise, with improved MM into both the PSL and EX fibers and also better polarization control. For this, I want to make a few changes at the EX table. 

  1. Replace existing fiber collimator with one of the recently acquired F220-APC-1064 collimators.
    • This gives an output mode of diameter 2.4mm with a beam divergence angle of 0.032 degrees (all numbers theoretical - I will measure these eventually but we need a beam path of ~5m length in order to get a good measurement of this collimated beam).
    • I believe it will be easier to achieve good mode matching into this mode rather than with the existing collimator. 
    • Unfortunately, the mount is still going to be K6X and not K6XS. 
  2. Improve mode-matching into fiber.
    • I used my measurement of the Innolight NPRO mode from 2016, a list of available lenses, and some measured distances to calculate a solution that gives theoretical 100% overlap with the collimator mode, that has beam diameter 2.4mm, located 80cm from the NPRO shutter head location (see Attachment #1).
    • The required movement of components is schematically illustrated in Attachment #2.
    • One of the required lens positions is close to the bracket holding the enclosure to the table, but I think the solution is still workable (the table is pretty crowded so I didn't bother too much with trying to find alternative solutions as all of them are likely to require optics placed close to existing ones and I'd like to avoid messing with the main green beam paths.
    • I will attempt to implement this and see how much mode matching we actually end up getting.
  3. Install a PBS + HWP combo in the EX fiber coupling path.
    • This is for better polarization control.
    • Also gives us more control over how much light is coupled into the fiber in a better way than with the ND filters in current path.
  4. Clean EX fiber tip.
  5. Dump a leakage IR beam from the harmonic separator post doubling oven, which is currently just hitting the enclosure. It looks pretty low power but I didn't measure it.
  6. Re-install EX power monitoring PD.

Barring objections, I will start working on these changes later today.

Attachment 1: EX_fiber_MM.pdf
EX_fiber_MM.pdf
Attachment 2: EX_fiber_changes.png
EX_fiber_changes.png
  13783   Tue Apr 24 10:10:43 2018 gautamUpdateComputer Scripts / ProgramsParticle swarm hyper parameter optimization

I'm copying and pasting Nikhil's email here as he was unable to login to the elog (but should now be able to in order to reply to any comments, and add more details about this test, motivation, methodology etc).

I did some post-processing after running the grid search. The following steps were carried out:

1) Selected those sets whose cost fun were less than a specific threshold (here 10000)

2) Next task was to see if the parameters of these good solutions had some pattern

3) I used a dimensionality reduction technique called t-SNE to project the 6 dimensional parameter space to 2 dim (for better visualization )

4) Made a scatter plot of these (see fig )

5) Used K-Means to find the clusters in this data

6) MarkerSize & Color reflect the cost fun. Bigger the marker size means better the solution.

7) Visual inspection implied cluster 5 had the best ranking points & more than any other cluster

8) These points had the following Parameter set: Workers {20,40}, SwarmSize {500}, MaxIter {500}, Self Adjustment {1}, Social Adjustment {1}, Tolerance {1e-3,1e-8} 

     See fig: for the box plot 

9) It looks like is a particular set of values rather than individual values that gives the best results.

 

Attachment 1: ClusterFminScaled.png
ClusterFminScaled.png
Attachment 2: ClusterID_5.png
ClusterID_5.png
  13782   Tue Apr 24 09:10:20 2018 KiraUpdatePEMfinal setup sketch

I've attached the final sketch for the panel on the granite block.

Attachment 1: EX-can-panel_1.pdf
EX-can-panel_1.pdf
  13781   Tue Apr 24 08:36:47 2018 johannesConfigurationGeneralAux Laser LD dying? (AS port laser injection)

In September 2017 I measured ~150mW output power, which was already kind of low. What are the chances of getting this one repaired? Steve, can you please check the serial number? It's probably too old like the other ones.

Quote:

I suspect that the LD of the aux laser is dying.
- The max power we obtain from this laser (700mW NPRO) is 33mW. Yes, 33mW. (See attachment 1)

 

  13780   Mon Apr 23 20:06:35 2018 ranaUpdatePEMPID tune

This shows a step response of the EX seis temp control with K_I = -1 and K_P = -0.1. The time constants for both heatup and cooldown are ~2 hours.

I'm not so sure if the PID code itself makes sense though:

  # The basic finite-difference PID approximation
  e[0] = (p-s)
  print("Error signal = {}" .format(e[0])) 

  # These are the main equations of the PID Process
  u[0] = u[1]
  u[0] = u[0] + Kp * (e[0] - e[1])
  u[0] = u[0] + Ki * (e[0])
  u[0] = u[0] + Kd * (e[0] - 2*e[1] + e[2])

 

Seems like the Proportional term uses the difference (or derivative) of the error signal. This makes it more likely to pick up some high frequency noise; maybe we should low pass this signal somewhat, or at least implement a running average.

Since we still don't have an out of loop sensor or a PSL room temperature monitor or a particle counter in the frames, I've disabled the PID loop to see how much the can temperature varies with no feedback. Please leave it this way for a few days.

Attachment 1: HeaterTest.png
HeaterTest.png
  13779   Sat Apr 21 20:25:12 2018 gautamUpdateALSPSL fiber pickoff status

Seems like there is still a bit of variation in the power in the two polarizations, though it is much smaller now, at the ~5% level (see Attachment #1). Since the pattern is repeating itself over the day timescale, I think this effect is not because of rotation of the output coupler in the mount, but is in fact a temperature driven waveplate effect because of imperfect alignment at the input coupler (which itself is locked down). I'm going to rotate the input coupler by 5 degrees (old = 110 degrees, new=115degrees) to see if the situation improves...


gautam Apr 24 2pm: Steve suggested confirming the correlation by hooking up the PSL table temperature sensor. This used to be logged but since the c1psl ADC card failure, has not been recorded. Assuming the sensor and preamp still work fine, we can use the PSL diagnostic Acromag (whose channels I have hijacked to monitor polarization stability already) to at least temporarily monitor the temperature inside the PSL enclosure. I am in need of a DB25 breakout board for this purpose which I am missing right now, as soon as I obtain one, I'll hook this up...

Attachment 1: PSL-beatMouthPickoff.png
PSL-beatMouthPickoff.png
  13778   Sat Apr 21 20:19:05 2018 gautamUpdateGeneralMegatron hard-rebooted

I found megatron in a similar state to that which nodus was in yesterday. Clued by the fact that MCautolocker wasn't executing the mc scripts (as was evident from looking at the wall StripTool trace), I tried ssh-ing into megatron, but was unable to (despite it being responsive to ping requests). So I went into the VEA and plugged in a monitor to megatron - saw nothing on it. With no soft reboot options available, I power cycled the machine via the front panel button. It came back up smoothly. I manually restarted the autolocker, FSSslow and EX thermal control processes (the former two with initctl, while the latter runs in a tmux session). Everything seems alright for now. Not sure how long megatron has been dead for.

  13777   Fri Apr 20 23:36:28 2018 KevinUpdatePEMSeismometer BLRMs

Steve secured the GPS time server in the rack above the AA board and removed the wooden block that it was resting on. The new rack is shown in attachment 1.

I then opened the AA board to see why the channels aren't working. Even though the board was powered and outputting 4.6 V, none of the chips were getting power. I must have shorted something while trying to diagnose this and the board is no longer powered either.

The schematic is given in D990147. The D68L8EX filter is bypassed on all the channels, as can be seen in attachment 3, so the board isn't really doing anything. Rana suggested that we could just bypass the whole circuit by wiring the IN channels directly to the OUT channels going to the ADC. I'll try that next for a single channel.

Attachment 1: front.jpg
front.jpg
Attachment 2: back.jpg
back.jpg
Attachment 3: detail.jpg
detail.jpg
  13776   Fri Apr 20 16:25:08 2018 SteveUpdateWiki ETMX table layout uploaded to wiki

ETMX table layout uploaded with beam paths to the wiki.   laugh

  13775   Fri Apr 20 16:22:32 2018 gautamUpdateGeneralNodus hard-rebooted

Aidan called saying nodus was down at ~345pm. I was able to access it at ~330pm. I couldn't ssh in from my machine or the control room ones. So I went to 1X7 and plugged in a monitor to nodus. It was totally unresponsive. Since the machine wasn't responding to ping either, I decided to hard-reboot it. Machine seemed to come back up smoothly. I had trouble getting the elog started - it wasn't clear to me that the web ports were closed by default, so even though the startELOGD.sh script was running fine, the 8080 port wasn't open to the outside world. Anyways, once I figured this out, I was able to start the elog. DokuWiki also seems to be up and running now... 

  13774   Fri Apr 20 15:07:45 2018 KiraUpdatePEMfinal setup sketch update

If we lay the cable along the floor then it should be around 6' to the current setup and about 20' to the actual seismometer.

Edit: 16 gauge wire should be good.

Quote:

since we're just going from the short rack (not the tall rack) to the seismometer, can't we use a cable shorter than 45' ?

 

  13773   Fri Apr 20 00:26:34 2018 gautamUpdateALSFibers switched out

Summary:

I think the dominant cause for the fact that we were seeing huge swing in the power coupled into the fiber was that the beam being sent in was in fact not linearly polarized, but elliptically polarized. I've rectified this with the help of a PBS. Fiber has been plugged into my polarization monitoring setup. Let's monitor for some long stretch and see if the situation has improved.

Details:

  • The new fiber mount I ordered, K6XS, arrived today. I like it - it has little keys with which all DoFs can be locked. Moreover, it is compatible with the fixed collimators which IMO is the easiest way to achieve good mode-matching into the fiber. It is basically a plug-and-play replacement for the mounts we were using. Anyways, we can evaluate the performance over the coming days.
  • I installed it on the PSL table (started work ~10pm, HEPA turned up to maximum, PSL shutter closed).
  • But even with the new rotational DoF locking feature, I saw that slight disturbances in the fiber caused wild fluctuations in my polarization monitoring setup PD outputs. This was a useful tool through the night of checking the polarization content in the two special axes - Aidan had suggested using a heat gun but shaking the fiber a bit works well too I think.
  • The PM980 fiber has an alignment key that is aligned with the slow axis of the fiber - so it is a useful alignment reference. But even by perturbing the roational alignment about the vertical by +/-15 degrees, I saw no improvement in this behavior. So I began to question my assumption that the input beam itself had clean polarization content.
  • Since my pickoff beam has gone through a QWP and two PBSs, I had assumed that the beam was linearly polarized.
  • But by putting a PBS just upstream of the input fiber coupler, I could see a beam at the S-port with an IR card (while I expected the beam to be P-polarized).
  • OK - so I decided to clean up the input polarization by leaving this PBS installed. With this modification to the setup, I found that me shaking the fiber around on the PSL table didn't affect the output polarization content nearly as dramatically as before!!yes
  • The state I am leaving it in tonight is such that there is ~100x the power in the P-polarization output monitor as the S-polarization (PER ~ 20dB). I didn't try and optimize this too much more for now, I want to observe some long term trend to see if the wild power fluctuations have been mitigated.
  • The output coupler is mounted on the inferior K6X mount, and so there is the possibility that some drift will be attributable to rotation of the output coupler in it's mount. Thermally driven length changes / time varying stresses in the fiber may also lead to some residual power fluctuations. But I don't expect this to be anywhere near the ~25% I reported in the previous elog.
  • The rejected beam from the PBS was measured to be ~300 uW. I haven't dumped this properly, to be done tomorrow.
  • HEPA turned back down to 30%, PSL enclosure closed up, PSL shutter re-opened ~0030am.
  • Note that the EX and EY fiber coupled beams are also likely subject to the same problem. We have to double check. I think it's better to have a PBS in front of the input fiber coupler as this also gives us control over the amount of light coupled into the fiber.

Power budget:

Power in Measured power (Ophir, filter OFF)
@Input coupler, before PBS 4.4 mW
P-pol content @ input coupler 4.06 mW
S-pol (rejected) from PBS 275 uW
@Output coupler 2.6 mW (MM ~65%)

 

  13772   Thu Apr 19 20:41:09 2018 KojiConfigurationGeneralAux Laser LD dying? (AS port laser injection)

I suspect that the LD of the aux laser is dying.
- The max power we obtain from this laser (700mW NPRO) is 33mW. Yes, 33mW. (See attachment 1)
- The intensity noise is likely to be relaxation oscillation and the frequency is so low as the pump power is low. When the ADJ is adjusted to 0, the peak moved even lower. (Attachment 2, compare purple and red)
- What the NE (noise eater) doing? Almost nothing. I suspect the ISS gain is too low because of the low output power. (Attachment 2, compare green and red)

Attachment 1: Aux_laser_adj_Pout.pdf
Aux_laser_adj_Pout.pdf
Attachment 2: Aux_laser_RIN.pdf
Aux_laser_RIN.pdf
  13771   Thu Apr 19 18:23:51 2018 KiraUpdatePEMfinal setup sketch update

since we're just going from the short rack (not the tall rack) to the seismometer, can't we use a cable shorter than 45' ?

Quote:

I've added in the dimensions to my sketch.

the panel should be completely replaced like I described. We don't want to try to squeeze it in artificially and torque the wires. It just needs to be separated from the slab by a few more cm.

  13770   Thu Apr 19 17:15:35 2018 gautamUpdateIOOMore IMC NBing

Summary:

Today, I repeated the coil driver noise measurement. Now, the coil driver noise curve in the noise budget plot (Attachment #1) is the actual measurement of all 12 coils (made with G=100 SR560). I am also attaching the raw voltage noise measurement (input terminated in 50ohms, Attachment #2). Note that POX11 spectrum has now been re-measured with analog de-whitening engaged on all 3 optics such that the DAC noise contribution should be negligible compared to coil driver noise in this configuration. The rows in Attachment #2 correspond to 800 Hz span (top) and full span (bottom) on the FFT analyzer.

Details:

The main difference between this measurement, and the one I did middle of last year (which agreed with the expectation from LISO modeling quite well) is that this measurement was done in-situ inside the eurocrate box while last year, I did everything on the electronics benches. So I claim that the whole mess of harmonics seen in the raw measurements are because of some electronics pickup near 1X6. But even disregarding the peaky features, the floor of ~30nV/rtHz is ~6x than what one would expect from LISO modeling (~5nV/rtHz). I confirmed by looking that the series resistance for all 3 MC optics is 430ohms. I also did the measurement with the nominal bias voltages applied to the four channels (these come in via the slow ADCs). But these paths are low-passed by an 8th order low pass with corner @ 1Hz, so at 100 Hz, even 1uV/rtHz should be totally insignificant. I suppose I could measure (with EPICS sine waves) this low-pass filtering, but it's hard to imagine this being the problem. At the very least, I think we should get rid of the x3 gain on the MC2 coil driver de-whitening board (and also on MC1 and MC3 if they also have the x3 factor).

Attachment 1: IMC_NB_20180419.pdf
IMC_NB_20180419.pdf
Attachment 2: IMC_coilDriverNoises_20180419.pdf
IMC_coilDriverNoises_20180419.pdf
  13769   Thu Apr 19 12:23:30 2018 KiraUpdatePEMfinal setup sketch update

I've added in the dimensions to my sketch.

It seems like placing the two connectors right next to each other would allow both cables to just barely go through the hole in the block.

Quote:

Can you please add dimensions to the drawing, so we can see if things fit and what the cable lenghts need to be?

For the panel on the granite slab, we should use a thinner piece of metal and mount it with an offset so that the D-sub cable can be fished through the hole in the slab. The hole is wide enough for 2 cables, but not 2 connectors.

 

Attachment 1: heater_1_new.png
heater_1_new.png
Attachment 2: heater_2_new.png
heater_2_new.png
  13768   Thu Apr 19 11:29:11 2018 ranaUpdatePEMPID tune

Yesterday, I changed the P gain of the PID loop from zero to  +0.1. Seems good so far; will monitor for a couple days to see if we're in the right ballpark. Main issue in the stability may now be that the quantization noise is too big for the temperature sensor. If so, we should consider subtracting off the DC value (with a V ref) and then amplifying before ADC.

Attachment 1: Screen_Shot_2018-04-19_at_11.27.08_AM.png
Screen_Shot_2018-04-19_at_11.27.08_AM.png
  13767   Thu Apr 19 09:57:03 2018 gautamUpdateWikiAP and ETMX tables uploaded to wiki

The most up to date pictures of the AP table and ETMX table that Steve took have been uploaded to the relevant page on the wiki. It seems like the wiki doesn't display previews of jpg images - by using png, I was able to get the thumbnail of the attachment to show up. It would be nice to add beam paths to these two images. The older versions of these photos were moved to the archive section on the same page.

  13766   Thu Apr 19 01:04:00 2018 gautamConfigurationGeneralAS port laser injection

For the arm cavity ringdowns, I guess we don't need AS55/AS110 (although I think the camera will still be useful for alignment). But for something like RC Gouy phase characterization, I'd imagine we need the AS detectors to lock various cavities. So I think we should go for a solution that doesn't disturb the AS PD beams. 

It's hard to tell from the plot in the manual (pg 52) what exactly the relaxation oscillation frequency is, but I think it's closer to 600 kHz (is this characteristic of NdYAG NPROs)??  Is the high RIN on the light straight out of the NPRO? 

Quote:

We could use a simple mirror if we don't need AS110 and AS55, we could use a polarizing BS and work with s polarization, or we find a Faraday Isolator.


There is a strong oscillation around 210 kHz. The oscillation frequency decreases when the output power is turned down, the noise eater has no effect. 

  13765   Thu Apr 19 00:03:51 2018 gautamUpdateIOOMore IMC NBing

Summary:

As shown in the Attachments, it seems like IMC DAC and coil driver noise is the dominant noise source above 30Hz. If we assume the region around the bounce peak is real motion of the stack (to be confirmed with accelerometer data soon), this NB is starting to add up. Much checking to be done, and I'd also like to get a cleaner measurement of coil driver and DAC noise for all 3 optics, as there seems to be a factor of ~5 disagreement between the MC3 coil driver noise measurement and my previous foray into this subject. The measurement needs to be refined a little, but I think the conclusion holds.

Details:

  1. I had a measurement of the MC3 coil driver noise from ~2weeks ago when I was last working on this that I had not yet added to the NB.
  2. Today I added it. To convert from measured voltage noise to frequency noise, I assumed the usual 0.016N/A per coil number, which is probably a large source of systematic error.
  3. I define the "nominal" IMC operating condition as MC1 and MC3 having the analog de-whitening filters switched on, but MC2 switched off.
  4. So length noise should be dominated by coil driver noise on MC1 and MC3, and DAC noise on MC2.
  5. The measurement I had was made with the input to the coil driver board terminated in 50ohms. Measurement was made in-situ. The measurement has a whole bunch of 60Hz harmonics (despite the Prologix box being powered by a linear power adapter, but perhaps there are other ground loops which are coupling into the measurement). So I'd like to get a cleaner measurement tmrw.
  6. To confirm, Koji suggested some On/Off test by driving some broadband noise in the coils. I figured toggling the analog de-whitening, such that the DAC noise or coil driver electronics dominate is an equally good test.
  7. Attachment #2 shows the effect in arm error and control signal spectra. Note that I engaged analog de-whitening on all 3 optics for the red curves in this plot. But even leaving MC2 de-whitening off, I could see the read curve was below the black reference trace, which was taken with de-whitening off on all 3 optics.

Remarks:

Since I sunk some time into it already, the motivation behind this work is just to try and make the IMC noise budget add up. It is not directly related to lowering the IR ALS noise, but if it is true that we are dominated by coil driver noise, we may want to consider modifying the MC coil driver electronics along with the ITM and ETMs.

Attachment 1: IMC_NB_20180409.pdf
IMC_NB_20180409.pdf
Attachment 2: IMC_coils_20180418.pdf
IMC_coils_20180418.pdf
  13764   Wed Apr 18 22:46:23 2018 johannesConfigurationGeneralAS port laser injection

Using Gautam's Finesse file and the cad files for the 40m optical setup I propagated the arm mode out of the AS port. For the location of the 3.04 mm waist I used the average distance to the ITMs, which is 11.321 m from the beam spot on the 2 inch mirror on the AS table close to the viewport. The 2inch lens focuses the IFO mode to a 82.6 μm waist at a distance of 81 cm, which is what we have to match the aux laser fiber output to.

I profiled the fiber output and obtained a waist of 289.4 μm at a distance of 93.3 cm from the front edge of the base of the fiber mount. Next step is to figure out the lens placement and how to merge the beam paths. We could use a simple mirror if we don't need AS110 and AS55, we could use a polarizing BS and work with s polarization, or we find a Faraday Isolator.


While doing a beam scan with the razor blade method I noticed that the aux laser has significant intensity noise. This is seen on the New Focus 1611 that is used for the beat signal between PSL and aux laser, as well as on the fiber output PD. There is a strong oscillation around 210 kHz. The oscillation frequency decreases when the output power is turned down, the noise eater has no effect. Koji suggested it could be light scattering back into the laser because I couldn't find a usable Faraday Isolator back when I installed the aux laser in the PSL enclosure. I'll have to investigate this a little further, look at the spectrum, etc. This intensity noise will appear as amplitude noise of the beat note, which worries me a little.

power_out_fluctuation_DC.png      power_out_fluctuation_AC_zoom.png

Attachment 1: ASpath.svg.png
ASpath.svg.png
  13763   Wed Apr 18 20:33:19 2018 KevinUpdateGeneralseismometer interfaces

Steve, the pictures you posted are not the AA board I was referring to. The attached pictures show the board which is sitting beneath the GPS time server.

Attachment 1: front.jpg
front.jpg
Attachment 2: back.jpg
back.jpg
Attachment 3: connectors.jpg
connectors.jpg
  13762   Wed Apr 18 19:02:15 2018 gautamUpdateIOOMC spot centering scripts

I'm working on fixing these (and the associated MEDM scripts) up so that we can get some reliable readback on how well centered the spots are on the MC mirrors. Seems like a bunch of MEDM display paths were broken, and it looks like the optimal demod phases (to put most of the output in I quadrature) are not what the current iteration of the scripts set them to be. It may well be that the gains that convert demodulated counts to mm of spot offset are also not correct anymore. I ran the script ~4times in ~1 hour time span, and got wildly different answers for the spot centering each time, so I wouldn't trust any of those numbers atm...

As you can see in Attachment #1, I stepped the demod phase of one of the servos from -180 to 180 degrees in 5degree increments. The previously used value of 57degrees is actually close to the worst possible point (if you want the signal in the I quadrature, which is what the scripts assume).

I used Attachment #2 to change up the demod phases to maximize the I signal. I chose the demod phase such that it preserved the sign of the demodulated signal (relative to the old demod phases). I also made some StripTool templates for these, and they are in the MC directory. Doing the spot centering measurement with the updated demod phases yields the following output from the script:

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[0.72506586251144134, 7.1956908534034403, 0.54754354165324848, -0.94113388241389784, -3.5325883025464959, -2.4934027726657142]

Seems totally unbelievable still that we are so far off center on MC1 yaw. Perhaps the gains and calibration to convert from counts to mm of spot offset need to be rechecked.

Attachment 1: MCass_demodPhase.png
MCass_demodPhase.png
Attachment 2: MCass_demodPhases.png
MCass_demodPhases.png
  13761   Wed Apr 18 17:15:35 2018 ranaConfigurationComputersNODUS: no xmgrace for dataviewer

Turns out, there is no RPM for XmGrace on Scientific Linux 7. Since this is the graphic output of dataviewer, we can't use dataviewer through X windows until this gets fixed. CDS is looking into a xmGrace replacement, but it would be better if we can hijack a alt RH repo to steal a temporary xmgrace RPM. KT has been pinged.

ELOG V3.1.3-